# Dual-Source SPIR over a noiseless MAC without Data Replication or Shared Randomness

### v1

@@ Abstract
Information-theoretically secure Symmetric Private Information Retrieval (SPIR) is known to be infeasible over noiseless channels with a single server. Known solutions to overcome this infeasibility involve additional resources such as database replication, shared randomness, or noisy channels. In this paper, we propose an alternative approach for achieving SPIR with information-theoretic security guarantees, without relying on shared randomness, noisy channels, or data replication. Specifically, we demonstrate that it is sufficient to use a noiseless binary adder multiple-access channel, where inputs are controlled by two non-colluding servers and the output is observed by the client, alongside a public noiseless communication channel between the client and the servers. Furthermore, in this setting, we characterize the optimal file rates, i.e., the file lengths normalized by the number of channel uses, that can be transferred.
@@ I Introduction
Consider a client who wishes to download a file from a server such that (i) their file selection is kept private from the server and (ii) the client does not learn any other information beyond the selected file. This setting is referred to as Symmetric Private Information Retrieval (SPIR). SPIR is also known as oblivious transfer [3, 4], which is anterior to SPIR and a fundamental cryptographic building block [5].
@@
Under information-theoretic security guarantees, which is the focus of this paper, it is well known, e.g., [6], that SPIR between a client and a single server is infeasible over a noiseless communication channel. To overcome this impossibility result, two approaches have previously been considered. The first approach consists in considering additional resources at the client and server in the form of correlated randomness, which could, for instance, be obtained through a noisy channel between the client and the server. Specifically, for some classes of noisy channels, SPIR is known to be feasible under information-theoretic security, e.g., [7, 6, 8, 9]. The second approach consists in replicating data in multiple servers and assuming that the servers share randomness, e.g., [10, 11, 12, 13]. Note that with this second approach, a necessary assumption is that only a strict subset of servers can collude against the client, otherwise SPIR is infeasible as the setting reduces to the case of SPIR between a client and a single server over a noiseless channel. This paper aims to demonstrate the feasibility of a third approach.
@@
In this paper, we propose to perform SPIR under information-theoretic security guarantees without shared randomness, a noisy channel, or data replication. Specifically, we consider SPIR between one client and two non-colluding servers, using a noiseless multiple-access channel, whose inputs are controlled by the servers and output is observed by the client, and a public communication channel between the client and the servers. The two servers have independent content so that no data is replicated. In this setting, the client aims to obtain one file from each server such that (i) the file selection remains private from the servers, (ii) the unselected files remain unknown to the receiver, and (iii) one server does not learn anything about the content of the other server. As formally described in Section II, in our setting, the servers and the client can communicate over a noiseless public channel and have access to a noiseless binary adder multiple-access channel, but no noisy channels nor pre-shared correlated randomness are available at the parties. While information-theoretically secure SPIR is impossible if the client engages in independent protocols with each of the servers, i.e., with one server at a time and without any intervention of the other server, we show that a multiuser protocol between the client and the two servers can enable information-theoretically secure SPIR with positive rates for both servers simultaneously by leveraging the binary adder multiple-access channel, instead of shared randomness, a noisy channel, or data replication. The main idea of the achievability scheme is to use the communication of one sender as noisy resource for the other sender and the receiver. Perhaps surprisingly, this can be accomplished without time sharing between the two servers, i.e., the client can retrieve a file from each server simultaneously. More specifically, we begin by designing an achievability scheme for the case where each server holds only two files, and then build on this scheme to handle the case where each server has multiple files. The main challenge is in the proof of the converse to establish the optimality of our approach. As a problem of independent interest, in the converse proof, we solve an optimization problem involving the distribution that maximizes the conditional entropy of two Bernoulli random variables given their sum. Our main results fully characterize the capacity region for this setting.
Related works: The crux of our approach is the use of a multiple-access channel. Similar to our study, other recent works have considered SPIR over multiple-access channels, including [14, 15, 16]. Compared to [14, 15], we focus on a noiseless, rather than noisy, multiple-access channel. Additionally, the privacy constraint in [15] concerns the privacy of the client selection with respect to individual servers, whereas we consider the privacy of the client selection with respect to both servers jointly. In contrast to [15, 16], we consider a setting with more than two files and without data replication. We also note that prior related work includes the study of PIR over a specific type of multiple-access channel, formed by the product of noisy point-to-point channels, in [17].
@@ II Problem Statement
Notation: For $a,b\in\mathbb{R}$, define $\llbracket a,b\rrbracket\triangleq[\lfloor a\rfloor,\lceil b\rceil]\cap\mathbb{N}$. For any $z\in\{1,2\}$, define $\bar{z}\triangleq 3-z$. For a vector $X^{n}$ of length $n$, for any set $\mathcal{S}\subseteq\llbracket 1,n\rrbracket$, let $X^{n}[\mathcal{S}]$ denote the components of $X^{n}$ whose indices are in $\mathcal{S}$. In the following, all logarithms are taken base two.
@@
Figure 1: Dual-source symmetric private information retrieval between one client and two servers who have access to a public channel and a binary adder MAC as described in (1). The file selection of the client is $(Z_{1},Z_{2})\in\mathcal{L}_{1}\times\mathcal{L}_{2}$, i.e., the client wish to obtain $(F_{1,Z_{1}},F_{2,Z_{2}})$ from the servers with the constraints that the client must not learn information about $(F_{1,\mathcal{L}_{1}\backslash\{{Z}_{1}\}},F_{2,\mathcal{L}_{2}\backslash\{{Z}_{2}\}})$, $(Z_{1},Z_{2})$ must remain private from both servers, and Server $1$ must not learn information about Server $2$’s content, and vice versa for Server $2$.
@@
and operates as follows from time $t=1$ to time $t=n$,
  • The servers send $((X_{1})_{t},(X_{2})_{t})\in\mathcal{X}_{1}\times\mathcal{X}_{2}$ over the binary adder MAC (1) and the client observes $Y_{t}\triangleq(X_{1})_{t}+(X_{2})_{t}$, where $(X_{1})_{t}$ and $(X_{2})_{t}$ are functions of $V_{1,t}\triangleq(F_{1,\mathcal{L}_{1}},U_{1},(A_{1,i})_{i\in\llbracket 1,t-1\rrbracket})$ and $V_{2,t}\triangleq(F_{2,\mathcal{L}_{2}},U_{2},(A_{2,i})_{i\in\llbracket 1,t-1\rrbracket})$, respectively;
  • Next, the servers and the client are allowed to communicate, possibly interactively in $r_{t}$ rounds of communication, over a noiseless channel. Specifically, for $j\in\llbracket 1,r_{t}\rrbracket$, Servers 1 and 2 form and publicly send the messages $M_{1,t}(j)$ and $M_{2,t}(j)$, respectively, which are functions of $(V_{1,t},(M_{0,1,t}(m))_{m\in\llbracket 1,j-1\rrbracket})$ and $(V_{2,t},(M_{0,2,t}(m))_{m\in\llbracket 1,j-1\rrbracket})$, respectively. The client forms and publicly sends the messages $M_{0,1,t}(j)$ and $M_{0,2,t}(j)$, which are functions of $(Z_{1},U_{0},(A_{1,i})_{i\in\llbracket 1,t-1\rrbracket},Y^{t},(M_{1,t}(m))_{m\in\llbracket 1,j\rrbracket})$ and $(Z_{2},U_{0},(A_{2,i})_{i\in\llbracket 1,t-1\rrbracket},Y^{t},(M_{2,t}(m))_{m\in\llbracket 1,j\rrbracket})$, respectively. Let $A_{1,t}\triangleq(M_{0,1,t}(j),M_{1,t}(j))_{j\in\llbracket 1,r_{t}\rrbracket}$ and $A_{2,t}\triangleq(M_{0,2,t}(j),M_{2,t}(j))_{j\in\llbracket 1,r_{t}\rrbracket}$ represent all the messages publicly exchanged between the $t$-th and $t+1$-th channel use.
Define the entire public communication by $\mathbf{A}\triangleq(\mathbf{A}_{1},\mathbf{A}_{2})$ with $\mathbf{A}_{1}\triangleq(A_{1,t})_{t\in\llbracket 1,n\rrbracket}$ and $\mathbf{A}_{2}\triangleq(A_{2,t})_{t\in\llbracket 1,n\rrbracket}$. Finally, the client forms $\hat{F}_{1,Z_{1}}$, an estimate of $F_{1,Z_{1}}$, from $(Y^{n},U_{0},\mathbf{A}_{1},Z_{1})$, and $\hat{F}_{2,Z_{2}}$, an estimate of $F_{2,Z_{2}}$, from $(Y^{n},U_{0},\mathbf{A}_{2},Z_{2})$.
@@
Equation (2) ensures that the client obtains the selected files. Equations (3) and (4) ensure the client’s privacy by keeping the file selection $(Z_{1},Z_{2})$ private from Server $1$ and Server $2$, respectively. Note that $(F_{i,\mathcal{L}_{i}},X_{i}^{n},U_{i},\mathbf{A})$ represents all the information available at Server $i\in\{1,2\}$ at the end of the protocol. Equation (5) (respectively Equation (6)) ensures Server $2$’s (respectively Server $1$’s) privacy with respect to Server $1$’s (respectively Server $2$’s). Equation (7) ensures the servers’ privacy by keeping all the non-selected files private from the client. Note that $(Z_{1},Z_{2},Y^{n},U_{0},\mathbf{A})$ represents all the information available at the client at the end of the protocol
Note that, if the servers are colluding, then no positive rates are achievable, as the setting reduces to an SPIR setting between one client and one server over a noiseless channel, for which it is known that information-theoretic security cannot be achieved, e.g., [11].
@@ III Main Results
Figure 2: The x-coordinate (resp. y-coordinate) of $\mathcal{C}_{\textup{SPIR}}$ is defined as the SPIR capacity between Server $1$ (resp. Server $2$) and client in the absence of Server $2$ (resp. Server $1$), i.e., $\mathcal{C}_{\textup{SPIR}}=(0,0)$. $\mathcal{C}_{\textup{SPIR}^{2}}(L_{1},L_{2})$ is the dual-source SPIR capacity region characterized in Theorem 1.
@@
$\mathcal{C}_{\textup{SPIR}^{2}}(L_{1},L_{2})=\left\{(R_{1},R_{2}):(L_{1}-1)R_{1}+(L_{2}-1)R_{2}\leqslant\frac{1}{2}\right\}.$
@@
Moreover, if $L_{1}=L=L_{2}$ and the files on both servers have the same size, i.e, $nR_{1}=nR=nR_{2}$, then

### v2

@@ Abstract
Information-theoretically secure Symmetric Private Information Retrieval (SPIR) from a single server is known to be infeasible over noiseless channels. Known solutions to overcome this infeasibility involve additional resources such as database replication, shared randomness, or noisy channels. In this paper, we propose an alternative approach for achieving SPIR with information-theoretic security guarantees, without relying on shared randomness, noisy channels, or data replication. Specifically, we demonstrate that it is sufficient to use a noiseless binary adder multiple-access channel, where inputs are controlled by two non-colluding servers and the output is observed by the client, alongside a public noiseless communication channel between the client and the servers. Furthermore, in this setting, we characterize the optimal file rates, i.e., the file lengths normalized by the number of channel uses, that can be transferred.
@@ I Introduction
Consider a client who wishes to download a file from a server such that (i) their file selection is kept private from the server and (ii) the client does not learn any other information beyond the selected file. This setting is referred to as Symmetric Private Information Retrieval (SPIR). SPIR is also known as oblivious transfer [rabin1981exchange, goldreich1987play], which is anterior to SPIR and a fundamental cryptographic building block [kilian1988founding].
@@
Under information-theoretic security guarantees, which is the focus of this paper, it is well known, e.g., [crepeau1997efficient], that SPIR between a client and a single server is infeasible over a noiseless communication channel. To overcome this impossibility result, two approaches have previously been considered. The first approach consists in considering additional resources at the client and server in the form of correlated randomness, which could, for instance, be obtained through a noisy channel between the client and the server. Specifically, for some classes of noisy channels, SPIR is known to be feasible under information-theoretic security, e.g., [crepeau1988achieving, crepeau1997efficient, nascimento2008oblivious, ahlswede2013oblivious, Shekofteh2025B]. The second approach consists in replicating data in multiple servers and assuming that the servers share randomness, e.g., [gertner2000protecting, sun2018capacity, chen2020capacity, wang2019symmetric]. Note that with this second approach, a necessary assumption is that only a strict subset of servers can collude against the client, otherwise SPIR is infeasible as the setting reduces to the case of SPIR between a client and a single server over a noiseless channel. This paper aims to demonstrate the feasibility of a third approach.
@@
In this paper, we propose to perform SPIR under information-theoretic security guarantees without shared randomness, a noisy channel, or data replication. Specifically, we consider SPIR between one client and two non-colluding servers, using a noiseless multiple-access channel, whose inputs are controlled by the servers and output is observed by the client, and a public communication channel between the client and the servers. The two servers have independent content so that no data is replicated. In this setting, the client aims to obtain one file from each server such that (i) the file selection remains private from the servers, (ii) the unselected files remain unknown to the receiver, and (iii) one server does not learn anything about the content of the other server. As formally described in Section II, in our setting, the servers and the client can communicate over a noiseless public channel and have access to a noiseless binary adder multiple-access channel, but neither noisy channels nor pre-shared correlated randomness are available at the parties. While information-theoretically secure SPIR is impossible if the client engages in independent protocols with each of the servers, i.e., with one server at a time and without any intervention of the other server, we show that a multiuser protocol between the client and the two servers can enable information-theoretically secure SPIR with positive rates for both servers simultaneously by leveraging the binary adder multiple-access channel, instead of shared randomness, a noisy channel, or data replication. The main idea of the achievability scheme is to use the communication of one sender as noisy resource for the other sender and the receiver. Perhaps surprisingly, this can be accomplished without time sharing between the two servers, i.e., the client can retrieve a file from each server simultaneously. More specifically, we begin by designing an achievability scheme for the case where each server holds only two files, and then build on this scheme to handle the case where each server has multiple files. A key challenge lies in the proof of the converse to establish the optimality of our approach. As a problem of independent interest, in the converse proof, we solve an optimization problem involving the distribution that maximizes the conditional entropy of two Bernoulli random variables given their sum. Our main results fully characterize the capacity region of this setting, defined as the set of all achievable rate pairs $(R_{1},R_{2})$, where $R_{i}$ denotes the retrieval rate from Server $i$, given by the ratio of the requested file size to the total number of channel uses. Unlike prior works on SPIR with replicated or coded data, which use the total download cost from the servers as the performance metric, our setting considers the number of channel uses of the binary adder multiple-access channel as the performance metric.
Note that a noiseless binary adder multiple access channel can be modeled by physical-layer schemes where simultaneous transmissions on a shared medium yield an output equal to the number of active users. Each user activity is represented as a binary input (0 if inactive, 1 if active) such that the receiver observes 0 when no one transmits, 2 when both transmit, and 1 when only one is active. In the last case, the receiver knows a single user sent a signal but cannot tell which one. This effect could be realized in several ways. For example, in simultaneous wireless transmissions with energy detection, each transmitter either emits a short signal or remains silent on the same frequency, allowing an energy detector at the receiver to determine how many transmitters were active. In power line communication backscatter, devices modulate the reflection of an existing signal on the shared wiring so that the combined reflections reveal the number of active devices. In optical channels, each transmitter could emit light if active, and the signals can be combined at a photodetector so that the measured intensity indicates how many transmitters were active. Realizing this effect in practice requires at least transmitter synchronization, controlled signal amplitudes, and limited channel variability to preserve linear superposition.
Related works: The crux of our approach is the use of a noiseless multiple-access channel. By contrast, other recent works have considered SPIR over noisy multiple-access channels, including [shekofteh2024spir, aghaee2025oblivious, elimelech2025spir]. Compared with [aghaee2025oblivious], where the privacy constraint concerns the client selection with respect to individual servers, our work considers privacy of the client selection with respect to both servers jointly. Moreover, while [aghaee2025oblivious] focuses on two files, we consider an arbitrary number of files. In [elimelech2025spir], the setting involves an arbitrary number of non-communicating and replicated servers over a Gaussian multiple-access channel, and joint channel-SPIR schemes that do not require common randomness among the servers are proposed. Our setting also does not rely on common randomness, but unlike [elimelech2025spir], it avoids data replication and instead considers two servers with independent content. In [shekofteh2024spir], for a specific class of noisy multiple-access channels, positive rates are achievable even when all servers collude, in the scenario where the file selection index is identical across servers, as may happen when files are distributed across multiple servers.
Similar to our setting, a noiseless binary adder multiple-access channel was also studied in [pei2023capacity]. The main difference is that [pei2023capacity] considers replicated servers, whereas we do not allow data replication. Another difference is that [pei2023capacity] focuses on the case of two files, whereas we consider a setting with an arbitrary number of files. We also note that prior related work includes the study of PIR with non-communicating, replicated servers over a multiple-access channels formed by the product of noisy point-to-point channels [banawan2019noisy], as well as over a Gaussian multiple-access channel in [shmuel2021private] and in [elimelech2025pir], in the case where server responses are public.
@@ II Problem Statement
Notation: For $a,b\in\mathbb{R}$, define $\llbracket a,b\rrbracket\triangleq[\lfloor a\rfloor,\lceil b\rceil]\cap\mathbb{N}$. For any $z\in\{1,2\}$, define $\bar{z}\triangleq 3-z$. For a vector $X^{n}$ of length $n$, for any set $\mathcal{S}\subseteq\llbracket 1,n\rrbracket$, let $X^{n}[\mathcal{S}]$ denote the components of $X^{n}$ whose indices are in $\mathcal{S}$. Let $\oplus$ denote addition modulo two. In the following, all logarithms are taken base two. We also give a notation overview in Table I.
TABLE I: Notation overview.
Notation Description
$\llbracket a,b\rrbracket$ Integer interval $\llbracket a,b\rrbracket\triangleq[\lfloor a\rfloor,\lceil b\rceil]\cap\mathbb{N}$
$\bar{z}$ For $z\in\{1,2\}$, $\bar{z}\triangleq 3-z$
$\oplus$ Addition modulo 2
$n$ Number of channel uses
$(X_{i})_{t}$ Channel input of Server $i\in\{1,2\}$ at time $t\in\llbracket 1,n\rrbracket$
$Y_{t}$ Output of the binary adder MAC at time $t\in\llbracket 1,n\rrbracket$, $Y_{t}\triangleq(X_{1})_{t}+(X_{2})_{t}$
$Y^{t}$ Sequence of outputs up to time $t\in\llbracket 1,n\rrbracket$, $Y^{t}\triangleq(Y_{i})_{i\in\llbracket 1,t\rrbracket}$
$X_{i}^{n}[\mathcal{S}]$ Components of vector $X_{i}^{n}\triangleq((X_{i})_{1},\dots,(X_{i})_{n})$ indexed by $\mathcal{S}\subseteq\llbracket 1,n\rrbracket$, $i\in\{1,2\}$
$L_{i}$ Number of files stored at Server $i\in\{1,2\}$
$\mathcal{L}_{i}$ Set of file indices for Server $i\in\{1,2\}$, $\mathcal{L}_{i}\triangleq\llbracket 1,L_{i}\rrbracket$, $i\in\{1,2\}$
$F_{i,l}$ File $l\in\mathcal{L}_{i}$ stored at Server $i\in\{1,2\}$, uniformly distributed over $\{0,1\}^{nR_{i}}$
$U_{0},U_{1},U_{2}$ Local randomness available at the client, Server 1, and Server 2, respectively
$Z_{1},Z_{2}$ Client’s file selection for Server 1 and Server 2, respectively
$\mathbf{A}$ Total public communication
$\hat{F}_{i,Z_{i}}$ Client’s estimate of the requested file $F_{i,Z_{i}}$, $i\in\{1,2\}$
@@
[…]
@@
and operates as follows,
  • For time $t=1$ to time $t=n$, the servers send $((X_{1})_{t},(X_{2})_{t})\in\mathcal{X}_{1}\times\mathcal{X}_{2}$ over the binary adder MAC (1) and the client observes $Y_{t}\triangleq(X_{1})_{t}+(X_{2})_{t}$, where $(X_{1})_{t}$ and $(X_{2})_{t}$ are functions of $(F_{1,\mathcal{L}_{1}},U_{1})$ and $(F_{2,\mathcal{L}_{2}},U_{2})$, respectively. Define $Y^{n}\triangleq(Y_{t})_{t\in\llbracket 1,n\rrbracket}$.
  • Next, the servers and the client are allowed to communicate, possibly interactively in $r$ rounds of communication, over a noiseless channel. Specifically, for $j\in\llbracket 1,r\rrbracket$, Servers 1 and 2 form and publicly send the messages $M_{1}(j)$ and $M_{2}(j)$, respectively, which are functions of $(F_{1,\mathcal{L}_{1}},U_{1},(M_{0,1}(m))_{m\in\llbracket 1,j-1\rrbracket})$ and $(F_{2,\mathcal{L}_{2}},U_{2},(M_{0,2}(m))_{m\in\llbracket 1,j-1\rrbracket})$, respectively. The client forms and publicly sends the messages $M_{0,1}(j)$ and $M_{0,2}(j)$, which are functions of $(Z_{1},U_{0},Y^{n},(M_{1}(m))_{m\in\llbracket 1,j\rrbracket})$ and $(Z_{2},U_{0},Y^{n},(M_{2}(m))_{m\in\llbracket 1,j\rrbracket})$, respectively.
Define $\mathbf{A}_{1}\triangleq(M_{0,1}(j),M_{1}(j))_{j\in\llbracket 1,r\rrbracket}$ and $\mathbf{A}_{2}\triangleq(M_{0,2}(j),M_{2}(j))_{j\in\llbracket 1,r\rrbracket}$. The entire public communication is denoted by $\mathbf{A}\triangleq(\mathbf{A}_{1},\mathbf{A}_{2})$. Finally, the client forms $\hat{F}_{1,Z_{1}}$, an estimate of $F_{1,Z_{1}}$, from $(Y^{n},U_{0},\mathbf{A}_{1},Z_{1})$, and $\hat{F}_{2,Z_{2}}$, an estimate of $F_{2,Z_{2}}$, from $(Y^{n},U_{0},\mathbf{A}_{2},Z_{2})$.
@@
$\lim_{n\to\infty}\mathbb{P}\left[(\widehat{F}_{1,Z_{1}},\widehat{F}_{2,Z_{2}})\neq(F_{1,Z_{1}},F_{2,Z_{2}})\right]=0$ (2)
$\hskip 120.00018pt\text{(File recovery at Client)},$
$\lim_{n\to\infty}I\left(F_{1,\mathcal{L}_{1}}X_{1}^{n}U_{1}\mathbf{A};Z_{1}Z_{2}\right)=0$ (3)
$\hskip 120.00018pt\text{(Client privacy w.r.t. Server 1)},$
$\lim_{n\to\infty}I\left(F_{2,\mathcal{L}_{2}}X_{2}^{n}U_{2}\mathbf{A};Z_{1}Z_{2}\right)=0$ (4)
$\hskip 120.00018pt\text{(Client privacy w.r.t. Server 2)},$
$\lim_{n\to\infty}I(F_{1,\mathcal{L}_{1}}X^{n}_{1}U_{1}\mathbf{A};F_{2,\mathcal{L}_{2}})=0$ (5)
$\hskip 110.00017pt\text{(Server 2 privacy w.r.t. Server 1)},$
$\lim_{n\to\infty}I(F_{2,\mathcal{L}_{2}}X^{n}_{2}U_{2}\mathbf{A};F_{1,\mathcal{L}_{1}})=0$ (6)
$\hskip 110.00017pt\text{(Server 1 privacy w.r.t. Server 2)},$
@@
Equation (2) ensures that the client obtains the selected files. Equations (3) and (4) ensure the client’s privacy by keeping the file selection $(Z_{1},Z_{2})$ private from Server $1$ and Server $2$, respectively. Note that $(F_{i,\mathcal{L}_{i}},X_{i}^{n},U_{i},\mathbf{A})$ represents all the information available at Server $i\in\{1,2\}$ at the end of the protocol. Equation (5) (respectively Equation (6)) ensures Server $2$’s (respectively Server $1$’s) privacy with respect to Server $1$’s (respectively Server $2$’s). Equation (7) ensures the servers’ privacy by keeping all the non-selected files private from the client. Note that $(Z_{1},Z_{2},Y^{n},U_{0},\mathbf{A})$ represents all the information available to the client at the end of the protocol.
Note that, if the servers are colluding, then no positive rates are achievable, as the setting reduces to an SPIR setting between one client and one server over a noiseless channel, for which it is known that information-theoretic security cannot be achieved, e.g., [sun2018capacity].
@@ III Main Results
Figure 2: The x-coordinate (resp. y-coordinate) of $\mathcal{C}_{\textup{SPIR}}$ is defined as the SPIR capacity between Server 11 (resp. Server 22) and client in the absence of Server 22 (resp. Server 11), i.e., $\mathcal{C}_{\textup{SPIR}}=(0,0)$. $\mathcal{C}_{\textup{SPIR}^{2}}(L_{1},L_{2})$ is the dual-source SPIR capacity region characterized in Theorem 1.
@@
$\mathcal{C}_{\textup{SPIR}^{2}}(L_{1},L_{2})\\ =\left\{(R_{1},R_{2}):(L_{1}-1)R_{1}+(L_{2}-1)R_{2}\leqslant\frac{1}{2}\right\}.$
@@
Moreover, if $L_{1}=L=L_{2}$ and the files on both servers have the same size, i.e., $nR_{1}=nR=nR_{2}$, then